OpenStudy (anonymous):

people please help me ! there are 4 people and 1.5 sandwiches . i want to give each person the same amount of sandwiches . how much does each person get?

OpenStudy (richyw):

\[\frac{3}{2} \frac{1}{4}=\frac{3}{8}\] each person gets \(\frac{3}{8}\) of a sandwich